we have a ping pong ball with r=3 cm and density = 85 kg/m3.calculate the force that acts on the ping pong ball and it is keeping that under water.density of water = 1000 kg/m3
1
Expert's answer
2017-03-02T15:27:06-0500
F=F_bouyancy-mg=(ρ_water-ρ_(p-p) )Vg=4/3 πr^3 (ρ_water-ρ_(p-p) )g
F=4/3 π〖(0.03)〗^3 (1000-85)9.8=1 N.